To 100 ml sample solution that contains K ion with unknown concentration is added a number of volume standard solution K ion 0,1 mol/L. The emf is measured at 298 K in the table. Calculate the concentration of K ion. (Note: Volume larutan standar = standard solution volume)

Solve this with the addition standard method.

Emf (mV) 136 150 160 166 172 176 No Volume larutan standar K* (mL) 0,00 1,00 2,00 3,00 4,00 5,00 2 4
Answer #1

In No. 2 : standard K ion concentration in solution (Cs) = 0.1 mol/L x 1 ml/100 ml = 0.001 mol/L

Emf for 1 = 136 mV

Emf for 2 = 150 mV

let Cx be the concentration of K ion in unknown solution

So,

Cx/(Cx + 0.001) = 136/150

150Cx = 136Cx + 0.136

Cx = 0.136/14 = 0.0097 mol/L

Concentratiion of K ion is solution = 0.0097 mol/L